OpenStudy (edwardcullen2017):

what is cubed to equal 125?

OpenStudy (anonymous):

5

OpenStudy (edwardcullen2017):

thanks awheeler :)

OpenStudy (anonymous):

5

hero (hero):

\[x^3 = 125\]\[\sqrt[3]{x} = \sqrt[3]{125}\]\[x = 5\]